OpenStudy (anonymous):

A 220-year old oak tree produced acorns at two different rates. Until it reached maturity, it produced an average of 900 acorns a year. Once it reached maturity, it produced an average of 2200 acorns a year. Angel Oak has produced about 289,000 acorns in its lifetime. At what age did it reach maturity?

OpenStudy (cbarredo1):

Let the age at maturity be M 900 M + ( 220 - M) 2200 = 289000 900M + 484000 - 2200M = 289000 1300M = 195000 M = 195000 /1300 = 150 YEARS ANSWER